OpenStudy (batman19991):

If 9/24 =3/x , then x is can someone help me with this please

OpenStudy (anonymous):

do you have a general idea?

OpenStudy (batman19991):

no i dont i dont know how to do it at all

OpenStudy (anonymous):

I would first put the equation in simplest form

OpenStudy (anonymous):

so 9/24 can be reduced to 3/8

OpenStudy (anonymous):

so 3/8=3/x

OpenStudy (anonymous):

do you have choices?

OpenStudy (batman19991):

no

OpenStudy (anonymous):

x=8

OpenStudy (anonymous):

because 3/8=3/x and to equal the left side it has to bee x=8

OpenStudy (batman19991):

how did you figure that out though

OpenStudy (anonymous):

9/24=3/x we want to reduce 9/24 to the simplest form so we can figure that out

OpenStudy (anonymous):

now 9/24 can be reduced to 3/8 so our new equation is 3/8=3/x

OpenStudy (batman19991):

oh ok thanks

OpenStudy (anonymous):

In order to make them equal we have to figure out what is missing 3/8=3/x and we can infer that x has to equal 8 to make it equal

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

just divide and you get 63 miles per hr